Grade 12Magical Mathematics[Interesting Approach]

Q.
5 boys and 3 girls are to be arranged around a circular table such that B1 and G1 do not sit together.
What I did was
Total ways to arrange - Total ways in which B1 and G1 are together.
i.e. 7!-6!(2!)
But the answer is 5!*6!.
What am I doing wrong?

Rhitvik Anand
7 Years agoGrade 12

1 Answer

Arun
7 Years ago

Rule for Circular Permutations -

(n - 1)! = Clockwise + Anticlockwise arrangement.

Ex. Seating arrangement of persons round a table. 

- wherein

DISTINCT arrangement.

 

 Number of ways = Total - when B1 and G1 sit together 

Total ways to seat 8 people on table = 7!

When B1 and G1 sit together =6!\times 2!

Number of ways =7!-2\times 6!=6!(7-2)=5\times6!